Homework 6 (CH25 and CH26) Problem 25.39 2 of8 A11.0nC charge is at z = 0ctn and a-1.2 nC charge is at z 4 cm
Answer #1

There are two points where net potential due to both charge suppose to be zero

Between the charges

Let us consider the zero point is in between the two charges at a distance d from 11 NC

So

k q1 / d = K q2 / (4 - d)

11/d = 1.2 / ( 4 - d)

d= 3.61 cm

To the right of - 1.2 nC

Again Let us consider that the point is at a distance d from 11 nC and distance of that point from - 1.2nC becomes (d - 4)

Now

k q1 / d = K q2 / ( d - 4)

11/ d = 1.2 / ( d - 4)

d = 4.49 cm

So 3.61 cm and 4.49 cm

Comment in case any doubt please rate my answer ..